A machine with a cost of $100,000 and an estimated useful life of 200,000 units with no salvage value, was used to produce 20,000 units the first year and 30,000 units the second year. Using the units-of-production method, depreciation for the first and second years would be:

A)$10,000 and $15,000

B)$40,000 and $60,000

C)$20,000 and $30,000

D)Cannot be determined from the data given.
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(100,000/200,000)=$0.5/unit

Hence depreciation for

First year=(0.5*20,000)=$10,000

Second year=(0.5*30,000)=$15,000

Hence the correct option is:

A)$10,000 and $15,000